A light beam is traveling from medium 1, where its speed is V1 to medium 2 where its speed is V2= 0.5v2. if the first medium has index of refraction n2= 1.45 what is the index of refraction of the second medium?
Answer #1

Given,

speed in medium 1= v1

index of refraction of medium 1, n1 = 1.45

speed in medium 2, v2 = 0.5*v1

index of refraction of medium 2, n2 = ?

According to Snell's law

n1*v1 = n2*v2

1.45*v1 = n2*0.5*v1

n2 = 2.9